# bibliography.bib

### v1

@@ IV-C Environment Distillation: JaxLrauv UTracking
LRAUV Sim simulates the dynamics of a LRAUV robot at the millisecond level. However, LRAUV vehicles take actions only every few seconds to minutes. Morover, the tracking missions we are interested in span several hours. For this reason, it would be beneficial to simulate the environment state at a much larger time scale. Formally, we are interested in modeling the new position in the 2D plane $\mathbf{p}_{t+1}\in\mathbb{R}^{2}$ after a delta time $\delta_{t}$ given the actual robot position $\mathbf{p}_{t}$, absolute speed $v$ and rudder angle $\gamma$. Instead of reinventing the wheel and re-building an accurate physical model which is already provided by Gazebo, we want to approximate the change in the robots’ heading $\delta_{\psi}$ in order to use a simplified trajectory model, i.e. $\mathbf{p}_{t+1}=\mathbf{p}_{t}+v\delta_{t}\begin{bmatrix}\cos(\psi_{t}+\delta_{\psi}),\sin(\psi_{t}+\delta_{\psi})\end{bmatrix}$. Fixing $v$ and $\delta_{t}$, we parametrize $\delta_{\gamma}$ in respect to the rudder angle $\theta(\gamma)$. To approximate $\theta(\gamma)$ we follow an iterative distillation process: we collect trajectories in PyLrauv, build a simplified model, train an agent on that, and collect new data using the trained agent to improve the model distillation. We found that using an ensemble of linear models it is possible to approximate $\theta(\gamma)$ with a mean absolute error below 0.015 radians and a global $R^{2}$ score of 0.99. Such a simple model fails to grasp long time-dependency, and diverges when used autoregressively to generate long trajectories. Nevertheless, it performs well for shorter trajectories and is extremely faster than, for example, a $LSTM$. To improve the robustness of the system, we introduce Gaussian noise with a standard deviation of 0.02 radians, forcing the trained agent to handle trajectory uncertainty. In addition, we equip the agents with recurrent mechanisms. We found that training in this way in the simplified environment, the learned policies are directly transferable to Gazebo.
@@
The rest of the environment, including measurement noise, the probability of communication loss, partial observability and collision handling, is implemented manually following the observed statistics from the collected data. The final simplified environment is implemented in JAX to enable an end-to-end GPU-based RL pipeline, and is called JaxLrauv UTracking. The speedups achieved in different configurations of agents and targets are presented in LABEL:tab:speedup. These speedups allow to train a single-target tracking model that can be directly deployed in LRAUV Sim in as little as 10 minutes.
@@
TABLE I: Speedups Obtained with the Distilled Environment.
@@
IV-D Vectorized Particle Filter
While PF represent a state-of-the-art method for underwater tracking, their computational expense renders them inefficient for training, particularly in multi-agent scenarios. To mitigate this, [ivanscience] suggests employing Least Squares (LS) during training and reserving PF for testing. However, LS is only stable for static or slow-moving targets, limiting its scalability for faster targets and potentially intensifying the inherent non-stationarity of MARL. To facilitate efficient training, we implement a PF in JAX, utilizing vmap for GPU parallelization. By nesting vmapped functions, during training we update all particles across targets, agents, and parallel environments in a single pass, reducing the speed loss to only 30% compared to LS. As demonstrated in Figure 3, training with LS fails for targets moving faster than one-third of the agent’s velocity, whereas PF maintains stability.
Figure 3: Using Particle Filter or Least Squares to train an agent to follow a target at different velocities.
IV-E TransfMAPPO
The complexity of training multi-agent system with RL typically scales with the number of agents involved. To address this, we introduce TransfMAPPO, inspired by TransfQMix [transfqmix], which maintains the core concepts of the latter but uses PPO as the RL backbone instead of DQN, since PPO is more natural for vectorized environments. In TransfMAPPO, we implement a Transformer Actor and Critic that process the features describing the observed entities as vertices of a coordination latent graph learned through self-attention. The resulting policy and Value function are therefore invariant from the number of entities observed, allowing curriculum learning procedure described in subsection IV-F. More formally, the Transformer Critic observes the state of the environment represented as $n$ entities each one described by $z$ features, which in our case are the position, velocity and angles of the agents and targets. The Transformer Agent also observes the $n$ entities, but from its own partial observability perspective, which is constraint by the acoustic communication protocols defined in section IV. An overview of TransfMAPPO can be seen in Figure 4. Following [gallici2024simplifyingdeeptemporaldifference], we found that normalizing the entire network with LayerNormalization greatly improves stability and performances.
@@ IV-E TransfMAPPO
IV-F Curriculum Learning
Taking advantage of our Transformer-based architecture, we employ curriculum learning to progressively learn Multi-Agent general policies starting from a Single-Agent policy (see Figure 5). We begin with a large pre-training phase in which we train a single agent to maximize the tracking reward described in subsection IV-A with respect to an erratic-moving fast target (0.6x of its speed). We train for $10^{9}$ timesteps, corresponding to 1.5 days of compute (that would take months in Gazebo). In this phase, we maintain a short horizon (128 steps) before resetting the environment, as it significantly aids learning. The resulting policy is able to track a fast target with an error below 35 meters, but it fails to generalize to longer horizons. We therefore fine-tune the policy in progressively longer episodes (256, 512, and 1024 steps) for $10^{7}$ timesteps, until we observe that the policy can perfectly track a fast target for more than 10k steps, corresponding to 3.4 days of real-time tracking.
From this horizon-invariant base model, we transition to a multi-agent phase, where we fine-tune the policy for cooperative tasks. In particular, we consider two scenarios: (1) $n$ agents tracking $n$ targets, where we use the follow reward, and (2) $n$ agents tracking one very fast target (0.8x of their velocity), maximizing the same tracking reward used in the single-agent phase. Given that these two scenarios pose different cooperation and credit assignment problems, we reset the Critic parameters and maintain separate sets of fine-tuned agents for the two multi-agent fine-tuning branches, resulting in a final multi-target follow model and a multi-robot tracking model. In both cases, we fine-tune the model to cooperate with progressively more agents (5 and 3 at maximum, respectively) for a maximum $20^{8}$ timesteps. We observe that both resulting models can cooperate with a variable number of agents. In this phase, we also note that maintaining long environment horizons negatively affects training, so we reduce the horizon back to 256 steps. Unfortunately, while the resulting multi-agent tracking model retains the horizon invariance of the pre-trained single-agent model, the multi-target follow model loses some invariance in comparison.
@@ V-A Multi-Robot Tracking
The effectiveness of our curriculum method is demonstrated in Figure 6, where we compare the training curve of fine-tuning the model for the task of tracking a very fast target (0.8x the agents’ velocity) with 3 agents against training from scratch on the same task. The superior performance of curriculum learning is also evident in the reduced tracking error. In the same plot, we also include the training curve of single-agent PPO, which achieves less than half the returns of the multi-agent system.
@@
Figure 7: Multi-Robot Tracking Evaluation. The highglighted area represents the values seen in training.
To validate our final multi-robot tracking model based on TransfMAPPO, we conduct a comprehensive validation analysis using JaxLrauv Figure 10. Every data point in the validation phase is averaged across 1000 episodes, each with a length of 256 steps, following a target moving at 0.5x the agents’ velocity (unless otherwise specified). Specifically, we test how the tracking error is affected under three conditions: (1) when the target moves faster, (2) when its motion becomes more unpredictable (defined by the frequency of direction changes), and (3) when the agents’ motion is perturbed (simulating external forces such as ocean currents). In the last sub-figure we illustrate how the multi-agent system’s performance scales naturally with the number of agents used to track the target. Notably, even using only 2 agents significantly enhances the system’s robustness compared to a single-agent approach. We also highlight that the tracking error decreases monotonically with the number of agents, with 3 agents representing the sweet spot, achieving negligible collision probability while maintaining optimal tracking performance.
@@ V-B Multi-Target Tracking
To further test the effectiveness of TransfMAPPO, we train it from scratch alongside MAPPO in a simpler configuration where targets move slower (0.3x the agents’s speed) and in a more predictable manner. In this case, TransfMAPPO and MAPPO achieve similar scores with 2 agents, but MAPPO’s performance drops significantly with 3 agents. Surprisingly, TransfMAPPO learns very effectively even from scratch in the 5-agent-5-target scenario, confirming the sample efficiency obtained by decoupling a cooperation policy using Transformers.
@@ V-C Results in Gazebo
    Tracking a slow target (0.3$\times$ the agent’s velocity) for 300 steps using a single robot trained for 10 minutes in JaxLrauv.
@@
    Tracking a fast target (0.6$\times$ the agent’s velocity) for 1000 steps (equivalent to over 8 hours of real-time tracking) using our horizon-invariant base model.
@@
    Tracking a very fast target (0.8$\times$ the agent’s velocity) for 1000 steps using a multi-robot tracking fine-tuned model with 3 agents.
@@
    Tracking 3 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
@@
    Tracking 5 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
For the multi-target experiments, we reduced the number of steps to 300 due to the computational burden of simulating multiple robots in Gazebo. The results of these experiments are presented in Table II. Our findings demonstrate that even with only 10 minutes of training, we were able to develop an effective model capable of maintaining target tracking without losing the target. The multi-robot tracking model achieved a promising average tracking error of 3 meters, even when following a very fast target. This performance highlights a clear advantage over single-agent trackers following slower targets. Notably, the average agent-target distance increased when using multiple robots, reflecting a broader distribution of robots around the target space. The multi-target models performed as expected, with a small 5% probability of losing a single landmark. However, the multi-agent models exhibited an increased probability of collisions. This issue is likely attributable to the fact that the models were trained to avoid collisions only every 30 seconds, which can result in intra-step collisions in Gazebo. This underscores a critical area for future work to enhance the systems’ safety before deploying them in real-world marine missions.
TABLE II: Experimental Results in Gazebo
Config Avg. Agent-Target Distance Av. Tracking Error Prob. Collision (%) Prob. Lost (%)
1A,1T (Slow, 10Mins Train) 59.10 ± 32.29 5.89 ± 3.7 0.00 0.00
1A,1T (Fast) 121.01 ± 62.11 20.33 ± 28.38 0.00 7.14
3A,1T (Very Fast) 217.64 ± 89.85 3.03 ± 6.17 15.32 0.00
3A,3T (Moderate) 45.30 ± 26.11 5.29 ± 4.88 5.00 5.00
5A,5T (Moderate) 49.24 ± 30.92 4.25 ± 4.57 10.00 5.26
@@ VI Conclusions
We presented a method for scaling up Multi-Agent Reinforcement Learning (MARL) in the context of underwater tracking, demonstrating how it is possible to efficiently train MARL models purely on GPUs and deploy them in high-fidelity simulators, where MARL training would otherwise be extremely expensive or infeasible. Although this method relied on manual adjustments during the environment distillation phase, our overall pipeline represents a promising direction for closing the sim-to-real gap in MARL. A natural next step for this work is to enhance the safety of the multi-agent system and test it in real-world sea missions.
Figure 10: Multi-agent episodes collected in Gazebo. (a): Single-agent tracking, target velocity 0.66× of agent velocity. Video. (b): Multi-agent tracking, target velocity 0.88x of agents velocity. Video.
Figure 11: Agents trajectory according to target speed. Notice the change in the curvature of the agent’s trajectory to optimally track targets moving at different speeds.
Figure 12: Coordination when tracking a very fast target. Notice how agents react to a rapid change of direction of the target by waiting each other in order to not loose communication and then rapidly moving together.
Figure 13: Coordination in multi-target setting. Notice how agents resolve the ”traffic” problem by making circles to wait for other agents’ passing, and one target is dynamically reassigned from agent 3 to agent 1.
@@ ACKNOWLEDGMENT
This work acknowledges the Spanish Ministerio de Ciencia, Innovación y Universidades (BITER-ECO: PID2020-114732RBC31). This work is part of DIGI4ECO, which has received funding from the European Union’s Horizon Europe programme under grant agreement No 101112883. I. M. received financial support from the MCIN/AEI/10.13039/501100011033 and FSE+, grant agreement No RYC2022-038056-I. M. G. was partially founded by the FPI-UPC Santander Scholarship FPI-UPC_93. This work also acknowledges the ’Severo Ochoa Centre of Excellence’ accreditation (CEX2019-000928-S) and the Barcelona Super Computer Center.
[…]

### v2

@@ IV-C JaxLrauv UTracking
LRAUV Sim simulates the dynamics of an LRAUV robot at the millisecond level. However, in practice, LRAUV vehicles take actions only every few seconds to minutes, while the tracking missions of interest span several hours. For this reason, it is beneficial to simulate the environment state at a much larger time scale. Formally, we are interested in modeling the new position in the 2D plane $\mathbf{p}_{t+1}\in\mathbb{R}^{2}$ after a time step $\delta_{t}$, given the current robot position $\mathbf{p}_{t}$, absolute speed $v$, and rudder angle $\gamma$. Instead of re-building an accurate physical model (already provided by Gazebo), we approximate the change in the robot’s heading, $\delta_{\psi}$, in order to use a simplified trajectory model: $\mathbf{p}_{t+1}=\mathbf{p}_{t}+v\delta_{t}\begin{bmatrix}\cos(\psi_{t}+\delta_{\psi}),\sin(\psi_{t}+\delta_{\psi})\end{bmatrix}$. Fixing $v$ and $\delta_{t}$, we parametrize $\delta_{\psi}$ with respect to the rudder angle through a function $\theta(\gamma)$. To approximate $\theta(\gamma)$, we collect trajectories in PyLrauv and train a supervised model that predicts $\delta_{\psi}$ from $\theta(\gamma)$. We found that using a linear model for each fixed combination of $v$ and $\delta_{t}$, $\theta(\gamma)$ can be approximated with a mean absolute error below 0.015 radians and a global $R^{2}$ score of 0.99. Based on this, we build an ensemble of linear models for different combinations of $v$ and $\delta_{t}$. While such a simple model cannot capture long-term dependencies and diverges when used autoregressively for long trajectories, it performs well for shorter trajectories and is much faster than, for example, an LSTM or a handwritten physical system. To improve robustness, we introduce Gaussian noise with a standard deviation of 0.02 radians, forcing the trained agent to handle trajectory uncertainty. In addition, we equip the agents with recurrent mechanisms. We found that training in this simplified environment leads to policies that transfer directly to Gazebo.
@@
The rest of the environment, including measurement noise, communication loss, partial observability, and collision handling, is implemented manually based on data collected in Gazebo. The final simplified environment is implemented in JAX, enabling an end-to-end GPU-based RL pipeline. To further improve training efficiency, we implement a Particle Filter in JAX, allowing the tracking phase to also benefit from GPU acceleration. By nesting vectorized functions, we update all particles across targets, agents, and parallel environments in a single pass, achieving both high speed and precise tracking during training. The same vectorized PF algorithm is also integrated into PyLrauv for consistency. The speedups achieved across different agent–target configurations are reported in LABEL:tab:speedup. These gains make it possible to train a single-target tracking model that can be directly deployed in LRAUV Sim in as little as 10 minutes.
@@
TABLE I: Speedups Obtained with the Accelerated Environment. Steps Per Second (SPS) and relative speedup.
@@
IV-D TransfMAPPO
The complexity of training multi-agent systems with RL typically scales with the number of agents involved. To address this, we introduce TransfMAPPO, inspired by TransfQMix [15], which preserves the core concepts of the latter but uses PPO instead of DQN as the RL backbone, since PPO is naturally suited for vectorized environments. To leverage Transformers, we formalize the problem as learning a latent coordination graph through self-attention, where only the vertices of the graph are known. Specifically, the $n$ vertices (or tokens) correspond to the agents and targets present in the environment, each represented by $z$ features such as position, velocity, and orientation. This representation is permutation-invariant, allowing the Transformer to learn a coordination graph that is also invariant to the number of agents and targets—much like an LLM can learn semantic attention patterns over sentences of varying lengths. We distinguish between two types of latent representations: (1) the local latent graphs which can be learned by each agent aggregating noisy and potentially incomplete information from onboard sensors and inter-agent communications, and (2) the global latent graph learnable using the absolute states of all agents and targets, which is available only during training in simulation. Accordingly, we use two Transformer networks Figure 3: the TransformerAgents, deployed in a decentralized fashion (one per vehicle), which learns the local coordination graph to control the agents using only their local observations; and the TransformerCritic, which learns a value function for the joint action–state pair using the global information, and is employed for centralized training. The use of Transformers in both actor and critic improves coordination representation and yields policies and value functions that are invariant to the number of entities. This property enables the curriculum learning procedure described in subsection IV-E.
Following [31], we found that normalizing the entire network with LayerNormalization greatly improves stability and performances.
@@ IV-D TransfMAPPO
IV-E Curriculum Learning
Taking advantage of our Transformer-based architecture, we employ curriculum learning to progressively learn Multi-Agent policies starting from a Single-Agent policy (see Figure 4). We begin with a large pre-training phase in which we train a single agent to maximize the tracking reward described in subsection IV-A with respect to an erratic-moving fast target (0.6x of its speed). We train for $1^{10}$ timesteps, corresponding to 1.5 days of compute (that would take months in Gazebo). In this phase, we maintain a short horizon (128 steps) before resetting the environment, as it significantly aids learning. The resulting policy is able to track a fast target with an error below 35m, but it fails to generalize to longer horizons. We therefore fine-tune the policy in progressively longer episodes (256, 512, and 1024 steps) for $1^{8}$ timesteps, until we observe that the policy can perfectly track a fast target for more than 10k steps, corresponding to more than 3 days of real-time tracking.
From this horizon-invariant base model, we transition to a multi-agent phase, where we fine-tune the policy for cooperative tasks. In particular, we consider two scenarios: (1) $n$ agents tracking $n$ targets, where we use the follow reward, and (2) $n$ agents tracking one very fast target (0.8x of their velocity), maximizing the same tracking reward used in the single-agent phase. Given that these two scenarios pose different cooperation and credit assignment problems, we reset the Critic parameters and maintain separate sets of fine-tuned agents for the two multi-agent fine-tuning branches, resulting in a final multi-target follow model and a multi-robot tracking model. In both cases, we fine-tune the model to cooperate with progressively more agents (5 and 3 at maximum, respectively) for a maximum $2^{9}$ timesteps.
@@ V-A Multi-Robot Tracking
In Figure 5, we show the training curve for the most complex multi-robot tracking task we consider, i.e., three agents tracking a very fast target (0.8× the agents’ velocity). Specifically, we present results for training standard MAPPO (with an RNN) and TransfMAPPO from scratch, as well as the result of fine-tuning TransfMAPPO on this task following the curriculum procedure described in subsection IV-E. Note that training Transformers from scratch on this task is not more effective than using an RNN; however, the benefits of curriculum learning are evident in terms of expected returns and reduced tracking error.
@@
Figure 6: Multi-Robot Tracking Evaluation. The highlighted area represents the values seen in training.
In Figure 6, we perform a thorough evaluation of the final multi-target follow model obtained through curriculum training. Each data point in the validation phase is averaged over 1000 episodes, each lasting 256 steps, with the target moving at 0.5× the agents’ velocity (unless otherwise specified). Specifically, we assess how the tracking error is affected under three conditions: (1) when the target moves faster, (2) when its motion becomes more unpredictable (defined by the frequency of direction changes), and (3) when the agents’ motion is perturbed (simulating external forces such as ocean currents). In the last sub-figure, we illustrate how the performance of the multi-agent system scales with the number of agents used to track the target. As expected, even using only two agents significantly improves the system’s robustness compared to a single-agent approach. Robustness increases proportionally with the number of agents, both in terms of resistance to variations in target velocity, target motion (i.e., the interval of target direction changes), and perturbation noise (additional noise added to agents’ trajectories). We also highlight that the tracking error decreases monotonically with the number of agents, with three agents representing a sweet spot: achieving negligible collision probability while maintaining optimal tracking performance.
@@ V-B Multi-Target Tracking
To further test the effectiveness of TransfMAPPO, we train it from scratch alongside MAPPO in a simpler configuration where targets move slower (0.3x the agents’s speed) and in a more predictable manner (Figure 8). In this case, TransfMAPPO and MAPPO achieve similar scores with 2 agents, but MAPPO’s performance drops significantly with 3 agents. Surprisingly, TransfMAPPO learns very effectively even from scratch in the 5-agent-5-target scenario, confirming the sample efficiency obtained by learning a coordination graph using Transformers.
@@ V-C Results in Gazebo
TABLE II: Evaluation in JaxLrauv and PyLrauv of the curriculum training policies (* except for first line which refers to 10 minutes training).
Configuration Avg. Agent-Target Distance Avg. Tracking Error Probability of Collision (%) Probability of Loosing a Target (%)
JaxLrauv PyLrauv JaxLrauv PyLrauv JaxLrauv PyLrauv JaxLrauv PyLrauv
1 Agent, 1 Target (Slow)* 55.20 ± 30.11 59.10 ± 32.29 5.12 ± 3.1 5.89 ± 3.7 0.00 0.00 0.00 0.00
1 Agent, 1 Target (Fast) 110.23 ± 58.90 121.01 ± 62.11 17.40 ± 21.5 20.33 ± 28.38 0.00 0.00 5.00 7.14
3 Agents, 1 Target (Very Fast) 200.11 ± 85.70 217.64 ± 89.85 2.65 ± 5.30 3.03 ± 6.17 0.0 15.32 0.00 0.00
3 Agents, 3 Targets (Moderate) 42.15 ± 25.00 45.30 ± 26.11 4.85 ± 4.20 5.29 ± 4.88 0.5 5.00 3.50 5.00
5 Agents, 5 Targets (Moderate) 46.12 ± 28.75 49.24 ± 30.92 3.80 ± 4.00 4.25 ± 4.57 2.1 10.00 4.10 5.26
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
For the multi-target experiments, we reduced the number of steps to 300 due to the computational burden of simulating multiple robots in Gazebo. The results of these experiments are presented in Table II. For completeness, we also include in this table the results obtained in JaxLrauv with the same configurations, so that it is possible to appreciate how the learned policy behaves in a very similar way in both simulators, resulting in comparable evaluation measures. Our findings demonstrate that even with only 10 minutes of training, we were able to develop an effective model capable of maintaining target tracking without losing the target. The multi-robot tracking model achieved a promising average tracking error of 3m, even when following a very fast target. This performance highlights a clear advantage over single-agent trackers following slower targets. Notably, the average agent-target distance increased when using multiple robots, reflecting a broader distribution of robots around the target space. The multi-target models performed as expected, with a small 5% probability of losing a single landmark. However, the multi-agent models exhibited an increased probability of collisions. This issue is likely attributable to the fact that the models were trained to avoid collisions only every 30 seconds, which can result in intra-step collisions in Gazebo. This underscores a critical area for future work to enhance the systems’ safety before deploying them in real-world marine missions. Finally, in Figures 9 to 12 we present some of the trajectories obtained during the final evaluation, together with zoomed-in views of the emergent coordination strategies learned.
@@ VI Conclusions
We presented a method for scaling up Multi-Agent Reinforcement Learning (MARL) in the context of underwater tracking, demonstrating how it is possible to efficiently train MARL models purely on GPUs and deploy them in high-fidelity simulators, where MARL training would otherwise be extremely expensive or infeasible. Our overall pipeline represents a promising approach for closing the sim-to-real gap in MARL. A natural next step for this work is to enhance the safety of the multi-agent system and test it in real-world sea missions.
Figure 9: (a): Single-agent tracking, target velocity 0.66× of agent velocity. Video. (b): Multi-agent tracking, target velocity 0.88x of agents’ velocity. Video.
Figure 10: Agents’ trajectory according to target speed. Notice the change in the curvature of the agent’s trajectory to optimally track targets moving at different speeds.
Figure 11: Coordination when tracking a very fast target. Notice how agents react to a rapid change of direction of the target by waiting for each other in order not to lose communication and then rapidly moving together.
Figure 12: Coordination in a multi-target setting. Notice how blue and orange agents resolve the ”traffic” problem by making circles to wait for other agents’ passing, and one target is dynamically reassigned from one agent to another.
@@ ACKNOWLEDGMENT
This work acknowledges the Spanish Ministerio de Ciencia, Innovación y Universidades (BITER-ECO: PID2020-114732RBC31). This work is part of DIGI4ECO, which has received funding from the European Union’s Horizon Europe programme under grant agreement No 101112883. I. M. received financial support from the MCIN/AEI/10.13039/501100011033 and FSE+, grant agreement No RYC2022-038056-I. M. G. was partially funded by the FPI-UPC Santander Scholarship FPI-UPC_93. This work also acknowledges the ‘Severo Ochoa Centre of Excellence’ accreditation (CEX2024-001494-S) from AEI 10.13039/501100011033 and the Barcelona Supercomputer Center.